OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 6 3 4 Line 2 x y –6 7 9 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many @Tina_Russo

OpenStudy (dtan5457):

Give me a sec, I think i almost got this.

OpenStudy (dtan5457):

got it, let me explain

OpenStudy (anonymous):

Np

OpenStudy (dtan5457):

D. infinitely many this is because they are actually the SAME line, with the same slope, intercepts. your first set of points are (-3,6) and (3,4) your 2nd set of points are (-6,7) and (9.2) you want to put these into point-slope equation, so you must have the slope y2-y1/x2-x1 do that for each and you realize both slopes are -1/3 then plug into point slope equation for first one I will use the points (-3,6) y-6=-1/3(x+3) y-6=-1/3x-1 y=-1/3x+5 2nd line i used the point (-6,7) y-7=-1/3(x+6) y-7=-1/3x-2 y=-1/3x+5 same lines, infinite solutions.

OpenStudy (dtan5457):

any questions don't hesitate to ask :)

OpenStudy (anonymous):

Thank You soo Much

OpenStudy (dtan5457):

np bud